# Measurement-Induced Multimode Squeezed Light Interferometers with Scalable Architectures

### v1

@@ Abstract
In the rapidly expanding realm of photonic quantum information processing, the role of multimode squeezed light interferometers is critical. These devices are pivotal in enabling a wide range of applications, from quantum computing and sensing to advanced communication systems. Although linear interferometers with full programmability have been realized with up to 32 modes [1, 2], scaling to hundreds of modes is hampered by exponential increases in optical losses with circuit depth. Here we present and experimentally demonstrate highly scalable, low-loss multimode squeezed light interferometer schemes, capitalizing on measurement-induced operations and a shallow circuit approach. We validate our approach through a 6-mode interferometer and subsequently demonstrate its scalability by extending the design to 400 modes. This development marks a step toward realizing large-scale Noisy Intermediate-Scale Quantum (NISQ) algorithm applications, employing shallow circuits and addressing the prevalent issue of optical losses in photonic quantum information processing.
@@ I Introduction
Multimode squeezed light interferometers, that is, multimode interferometers effecting a linear transformation on multiple input squeezed states, have become a cornerstone in the field of continuous variable photonic quantum information processing. They support a broad array of applications that extend across quantum computing, sensing, and communications. These interferometer systems are instrumental not just in generating entangled graph states [3, 4] but also play pivotal roles in advanced quantum neural networks and machine learning algorithms [5, 6], Gaussian Boson Sampling [7, 8, 9, 10], multipartite quantum key distribution [11, 12], and distributed quantum sensing [13, 14]. Moreover, they are key in realizing the ambitious goals of universal and fault-tolerant quantum computing through GKP state encoding, indicating their versatility and critical importance in the advancement of quantum technologies.
@@
Numerous realizations of multimode squeezed light interferometers have been achieved, employing various technological avenues such as free-space optics [11, 12, 14, 9], photonic integrated circuits [15, 16, 1], and fiber optics [17, 10, 9]. In some cases, the interferometer circuit is fully programmable while in others, the programmability is highly limited or even reduced to a single configuration. Notably, the realizations using photonic integrated circuits are often fully programmable, enabling an arbitrary unitary Gaussian transformation. However, despite the promise shown by scalable on-chip interferometers with full programmability, the ambitious goal of scaling up to hundreds of modes is severely constrained by the optical losses that are significantly higher than in free-space or fiber optics. It is thus unlikely that fully programmable interferometers can be scaled to a large mode number.
@@
Figure 1: Concept of measurement-induced squeezed light interferometers. a) Injecting $N$ single-mode squeezed states of light into a multimode interferometer with beam-splitters $\vec{T}$ and phase shifters $\vec{\phi}$, that implements an arbitrary linear transformation described by an $N\times N$ unitary matrix $U$, results in an entangled Gaussian state that can be described by a graph, b), where nodes and edges represent modes and correlations between them, respectively. c) A different way of obtaining the same graph state is to start from a larger graph state of $M$ modes and perform homodyne measurements on $M-N$ modes. The original correlations and the homodyne basis settings determine the resulting graph. d) The particular scheme we study in this paper consists of generating the initial graph state by injecting a sequence of squeezed states at time intervals $\tau$ into a two-mode, shallow-depth interferometer made up of unbalanced Mach-Zehnder interferometers with delay lengths of $\tau K_{i}$, for different integers $K_{i}$ – typically with $K_{1}=1$. This results in a multimode temporally encoded cluster state [18, 4], as illustrated by the graph in e). By partially measuring out this cluster state with variable homodyne basis settings $\vec{\theta}$, a reduced state corresponding to a graph as in b) is obtained. Note that the in-principle endless cluster state can be reduced to a finite size by $x$-basis measurements.
@@
While full programmability is preferred, it is not a mandatory prerequisite for efficiently exploring the solution space relevant to specific problems across all quantum algorithmic applications. For example, hybrid quantum-classical algorithms do not necessarily require the quantum devices to possess full universality. E.g., in the case of the variational quantum eigensolver, employed for determining the ground state energy of molecular systems, the parameterized circuit undergoes continuous adjustments based on the feedback from measurement outcomes until it reaches convergence. Provided that the circuit configuration resides within the viable solution space, there is no need for comprehensive programmability of the parameters, demonstrating a more flexible approach to quantum problem-solving.
@@
In this article, we introduce and experimentally validate a non-universal, but highly scalable and low-loss multimode squeezed light interferometer that hinges on the strategic use of measurement-induced operations to enable partial programmability. Harnessing a continuous-variable Gaussian multi-partite entangled state combined with interactions induced by high-efficiency homodyne measurements, we facilitate the implementation of an effective multimode squeezed light interferometer through an exceptionally shallow circuit architecture. We first demonstrate the basic principle by realizing a compact 6-mode squeezed light interferometer and demonstrating the tunability of such a system in producing a measurement-based graph state. Following this, we underscore the potential for scalability by successfully extending the interferometer to incorporate 400 squeezed modes. Our findings unlock the potential for large-scale implementations of NISQ algorithmic applications, including the promising domain of hybrid quantum-classical algorithms, which are now feasible with the utilization of shallow circuit designs.
@@ II Concept and theory
A conventional programmable squeezed light interferometer, shown in Fig. 1a, can generate arbitrary (zero-mean) Gaussian multi-mode states, mathematically represented as graphs where vertices correspond to optical modes and edges indicate correlations (Fig. 1b). While such an interferometer provides full programmability – allowing the construction of any desired graph through control of beam-splitting ratios, phases and input Gaussian states – the circuit depth increases proportionally with the number of modes, making it impractical to scale up. To address this limitation, we propose a fundamentally different approach: shifting the programmability from the interferometer itself to a measurement system. By first generating a fixed graph state of $M$ modes and then applying programmable quadrature measurements to $M-N$ modes (see Fig. 1c), a tailored graph with $N$ vertices is created.
@@
A specific class of initial $M$-mode graph states can be efficiently generated using a scalable platform based on time-domain correlations. A conceptual schematic of this graph state generator is depicted in Fig. 1d. The setup employs two sources of squeezed light directed through a cascaded set of interferometers with highly unbalanced arm lengths. The delays in the interferometers are integer multiples, denoted $K_{i},$ for $i=1,2,3...$, of the clock cycle $\tau$ of the squeezed light. The resulting states manifest correlations across two spatial output modes, A and B, as well as temporally among sequentially produced modes (Fig. 1e), with the complexity of these correlations increasing with the addition of more interferometers. As examples, by using a configuration of a single or two interferometers, it is possible to generate one- [18] and two-dimensional [4, 3] graph states with correlations suitable for continuous-variable quantum computing [19, 20].
@@
In the next phase, the programmability of the setup is realized through the homodyne measurement process applied to a subset of the modes. Initially, we may assume that these are all the temporal modes of spatial mode $A$, in which case $N=M/2$. Later, we discuss different measurement patterns. The quadrature angles of the homodyne detections are parameterized as $\vec{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{M-N})$. These measurements influence the entanglement structure of the remaining modes in spatial mode $B$, effectively tailoring the time-domain entangled state into a new $N$-mode graph state,
$|\Phi_{\text{out}}\rangle=\hat{L}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})\hat{S}(\vec{r}_{\text{eff}})|0\rangle^{\otimes N}$ (1)
@@
where $\hat{L}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})$ and $\hat{S}(\vec{r}_{\text{eff}})$ are the effective, measurement-induced operators corresponding to an $N$-mode linear optical network and squeezing of the $N$ input modes, respectively, and where the parameters $\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}},\vec{r}_{\text{eff}}$ are functions of the fixed circuit configuration $\{K_{i}\}$ and the programmable measurement angles $\vec{\theta}$ and (in principle, although not in our experiment) initial squeezing parameters $\vec{r}_{\text{in}}$. This measurement-induced programmability enables controllability of the properties of the output state while maintaining the scalability and low-loss benefits of the fixed interferometer architecture: The long delay lines in our system will introduce some amount of loss, but since they are only traversed once, these losses do not increase proportionally to the number of modes.
Thus, the resultant time-domain output of mode B mirrors the output of a traditional squeezed light interferometer, where correlations span different spatial outputs of a linear beam splitter network supplied with squeezed light as in Fig. 1a. To identify the effective circuit parameters for a given measurement-induced state, we first find – using the Williamson decomposition – the symplectic transformation that produces that state from vacuum. Next, we use the Bloch-Messiah-Euler decomposition to dissect the transformation into initial squeezing transformations $\hat{S}(\vec{r}_{\text{eff}})$ and passive unitary transformations $\hat{L}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})$.
Figure 2: Simulated relative deviation of measurement-induced unitaries from reference unitaries when employing the straight-forward “linear” measurement strategy on a 1D cluster state (a), vs. a “Knight’s distance” strategy (b) where a larger number of modes are used for inducing the circuit, thereby increasing the number of available control parameters. The panels in (c) show for interferometers of sizes $N=2$, $6$ and $10$ the deviations of a set of parameterized states from a random ensemble for the two different measurement strategies for different values of t and the (physical) input squeezing, $r_{\text{in}}$, used to generate the initial 1D cluster state. Since the squeezing parameters, $\vec{r}_{\text{test}}$, on the ensemble of random states is restricted and the deviations depend on the choice of such restrictions, they are best understood relative to the worst case scenario of zero input squeezing, $r_{\text{in}}=0$. It can be seen that the deviations increase overall as the size of the induced interferometers increases. However, for any $N$, as the input squeezing increases the deviations drop, with a tendency to saturate as $r\rightarrow\infty.$ Furthermore, zooming in on the difference between the linear and the Knight’s distance (KD) approach (d) shows an advantage of the latter which is magnified as the input squeezing increases. This advantage, however, diminishes as $N$ increases.
Our scheme is not universal; it is not able to emulate any arbitrary $N$-mode squeezed light interferometer. To gauge the scheme’s versatility, we turn to the concept of expressibility. The conventional definition of expressibility [21], that measures the distance between fidelity distributions of states obtained from a parameterized circuit and from a Haar-random ensemble, respectively, is only well-defined for finite-dimensional systems. We extend this idea to interferometers induced by measurements. Consider an $N$-mode squeezing parameter vector $\vec{r}_{\text{test}}$ and the corresponding “test” squeezed input states $\hat{S}(\vec{r}_{\text{test}})\left|0\right\rangle^{\otimes N}$ such that all $r_{i}$ are drawn randomly from a judiciously chosen distribution $p_{\text{test}}(r)$. N-mode unitary matrices $W,V$ and $X,Z$ are then drawn from the set of all linear passive optics unitary matrices ($\mathbb{H}$) and parametrized unitary matrices ($\mathbb{P}$), respectively. If the measures induced over the two sets of unitaries $\mathbb{H}$ and $\mathbb{P}$ are $\mu_{HH}$ and $\mu_{PP}$, respectively, and $\hat{U}_{W}$ is the operator corresponding to matrix $W$ (and similar for $V,X,Z$), we define the deviation of the parametrized states $\left|\phi_{X}\right\rangle=\hat{U}_{X}\hat{S}(\vec{r}_{\text{test}})\left|0\right\rangle^{\otimes N}$ from the states generated due to the random passive linear optics $\left|\psi_{W}\right\rangle=\hat{U}_{W}\hat{S}(\vec{r}_{\text{test}})\left|0\right\rangle^{\otimes N}$ using the operator [21]
[…]
The Hilbert-Schmidt norm of this operator is a measure of deviation and is evaluated in equation 3. Note that the first two terms are the $t^{th}$ moment of the intra set fidelities over the sets ($\mathbb{H}$) and ($\mathbb{P}$) and the last term is the $t^{th}$ moment of the inter-set fidelities. We estimate this quantity through sufficiently large sets of sampled matrices and use it as a measure of relative expressibility:
[…]

### v3

# Scalable architecture for measurement-induced squeezed light interferometers

@@ Abstract
Scalable interferometers lie at the heart of photonic quantum technologies, but their expansion has been fundamentally limited by optical losses that grow with circuit depth. Here, we introduce and experimentally demonstrate a measurement-induced architecture for multimode squeezed-light interferometers that overcomes this barrier. By shifting complexity from deep optical networks to programmable homodyne measurements, we realize effective transformations within a shallow, low-loss platform. We validate the principle with a six-mode device and extend it to a 400-mode interferometer, marking a leap in scale beyond conventional designs. Crucially, this strategy not only enables scalable squeezed light interferometry but also provides a powerful route to the generation of large-scale entangled states — a key requirement for quantum computing, simulation, and communication. Our results establish measurement-induced circuits as a practical pathway toward noisy intermediate-scale quantum (NISQ) applications, and future demonstrations of quantum advantage.
@@ I Introduction
Multimode squeezed light interferometers, that is, multimode interferometers effecting a linear transformation on multiple input squeezed states, have become a cornerstone in the field of continuous variable photonic quantum information processing. They support a broad array of applications that extend across quantum computing, sensing, and communications. Such interferometer systems are not only instrumental in generating large-scale entangled graph states [cluster19, GG1], but also underpin advanced architectures for quantum neural networks and machine learning algorithms [QNN1]. They form the core of Gaussian Boson Sampling [GBS1, GBS2, Thekkadath2022, USTC20, Xanadu21, JZ4], enabling both quantum advantage demonstrations and quantum simulation [Paesani2019, Zhu2024]. Beyond computing, they provide essential tools for distributed quantum sensing [DQS1, DQS2], and the preparation of complex non-Gaussian resource states [Su2019-ConversionGaussianStates, Eaton2022-MeasurementbasedGenerationPreservationa, Takase2023-GottesmanKitaevPreskillQubitSynthesizer, Larsen2025-IntegratedPhotonicSource]. Looking ahead, multimode squeezed light interferometers will be central to realizing the long-term vision of universal and fault-tolerant photonic quantum computing. In particular, when integrated with highly non-classical states such as Gottesman–Kitaev–Preskill (GKP) or cubic phase states, they offer the architectural backbone required to achieve fault tolerance and scalability.
@@
Numerous realizations of multimode squeezed light interferometers have been achieved, employing various technological avenues such as free-space optics [JZ4, DQS2, USTC20, 100Teleportation], photonic integrated circuits [Paesani2019, Larsen2025-IntegratedPhotonicSource], fiber optics [Larsen21, Xanadu21, USTC20], or even spectral mode mixing [Chen2014, Roslund2014, Cai2017]. In some cases, the interferometer circuit is fully programmable while in others, the programmability is highly limited or even reduced to a single configuration. Notably, the realizations using photonic integrated circuits are often fully programmable, enabling an arbitrary unitary Gaussian transformation. However, despite the promise shown by scalable on-chip interferometers with full programmability, the ambitious goal of scaling up to hundreds of modes is severely constrained by the optical losses that are significantly higher than in free-space or fibre optics. Furthermore, as the depth of the circuit increases the losses incurred also significantly increase. Thus, it is unlikely that fully programmable interferometers can be scaled to a large mode number.
@@
Figure 1: Concept of measurement-induced squeezed light interferometers. a) Injecting NN single-mode squeezed states of light into a multimode interferometer with beam-splitters $\vec{T}$ and phase shifters $\vec{\phi}$, that implements an arbitrary linear transformation described by an $N\times N$ unitary matrix $V$, results in an entangled Gaussian state that can be described by a graph, b), where nodes and edges represent modes and correlations between them, respectively. c) A different way of obtaining the same graph state is to start from a larger graph state of $M$ modes and perform homodyne measurements on MNM-N modes. The original correlations and the homodyne basis settings determine the resulting graph. d) The particular scheme we study in this paper consists of generating the initial graph state by injecting a sequence of squeezed states at time intervals $\tau$ into a two-mode, shallow-depth interferometer made up of unbalanced Mach-Zehnder interferometers with delay lengths of $\tau K_{i}$, for different integers $K_{i}$. This results in a multimode temporally encoded cluster state [Yokoyama13, cluster19], as illustrated by the graph in e). By partially measuring out this cluster state with variable homodyne basis settings $\vec{\theta}$, a reduced state corresponding to a graph as in b) is obtained. Note that the in-principle endless cluster state can be reduced to a finite size by $x$-basis measurements that remove correlations [Menicucci2011]. In our experiments, the input squeezing $\vec{r}_{\text{in}}$ is constant and $K_{1}=1$, $K_{2}=8$.
@@
While full programmability is preferred, it is not a mandatory prerequisite for efficiently exploring the solution space relevant to specific problems across all quantum algorithmic applications. For example, hybrid quantum-classical algorithms do not necessarily require the quantum devices to possess full universality. E.g., in the case of the variational quantum eigensolver [Peruzzo2014-VariationalEigenvalueSolver], employed for determining the ground state energy of molecular systems, the parametrized circuit undergoes continuous adjustments based on the feedback from measurement outcomes until it reaches convergence. Provided that the circuit configuration resides within the viable solution space, there is no need for comprehensive programmability of the parameters, demonstrating a more flexible approach to quantum problem-solving.
@@
In this article, we introduce and experimentally validate a non-universal, but highly scalable and low-loss multimode squeezed light interferometer that hinges on the strategic use of measurement-induced operations to enable partial programmability. Harnessing a continuous-variable Gaussian state entangled in multiple temporal modes combined with effective interactions induced by high-efficiency homodyne measurements, we facilitate the implementation of an effective multimode squeezed light interferometer through an exceptionally shallow circuit architecture. We first demonstrate the basic principle by realizing a compact 6-mode squeezed light interferometer and demonstrating the tunability of such a system in producing a measurement-based graph state. Following this, we underscore the potential for scalability by successfully extending the interferometer to incorporate 400 squeezed modes. Our findings unlock the potential for large-scale implementations of NISQ algorithmic applications, including the promising domain of hybrid quantum-classical algorithms, which are now feasible with the utilization of shallow circuit designs.
@@ II Concept and theory
A conventional programmable squeezed light interferometer, shown in Fig. 1a, is a multiport interferometer, described by an $N\times N$ unitary matrix $V$ and consisting of variable beamsplitters $\vec{T}$ and phase shifters $\vec{\phi}$, fed by squeezed vacuum input states $\hat{S}(\vec{r})\left|0\right\rangle^{\otimes N}$. This can generate arbitrary (pure and zero-mean) Gaussian multi-mode states $|\Phi_{\text{out}}\rangle=\hat{U}_{V}(\vec{T},\vec{\phi})\hat{S}(\vec{r})|0\rangle^{\otimes N}$, where $\hat{U}_{V}$ is the unitary operation of the interferometer $V$ on the $N$-mode Hilbert space. Mathematically, such pure states can also be represented as graphs [Menicucci2011-GraphicalCalculusGaussian] where vertices correspond to optical modes and edges indicate correlations (Fig. 1b). While such an interferometer provides full programmability – allowing the construction of any desired graph through control of beam-splitting ratios, phases and input Gaussian states – the circuit depth increases proportionally with the number of modes, making it impractical to scale up. To address this limitation, we propose a fundamentally different approach: shifting the programmability from the interferometer itself to a measurement system. By first generating an initial graph state of $M$ modes and then applying programmable quadrature measurements to MNM-N modes (see Fig. 1c), a tailored graph with $N$ vertices is created: The correlations between the remaining $N$ modes will change depending on their initial correlations with the MNM-N measured modes and on the choice of quadrature basis.
@@
A specific class of initial $M$-mode graph states can be efficiently generated using a scalable platform based on time-domain correlations [Menicucci2011]. A conceptual schematic of this graph state generator is depicted in Fig. 1d. The setup employs two sources of single-mode squeezed light with squeezing parameters $\vec{r}_{\text{in}}$ directed through a cascaded set of interferometers with highly unbalanced arm lengths. The delays in the interferometers are integer multiples, $K_{i}$, of the clock cycle $\tau$ of the squeezed light. The resulting states manifest correlations across two spatial output modes, A and B, as well as temporally among sequentially produced modes (Fig. 1e), with the complexity of these correlations increasing with the addition of more interferometers. As examples, by using a configuration of a single or two interferometers, it is possible to generate one- [Yokoyama13] and two-dimensional [cluster19, GG1] graph states with correlations suitable for continuous-variable quantum computing [Larsen21a, Larsen21b].
@@
In the next phase, the programmability of the setup is realized through the homodyne measurement process applied to a selected subset of modes. Initially, we consider the case where all temporal modes of spatial mode $A$ are measured, giving $N=M/2$. Alternative measurement patterns will be discussed later. The quadrature angles of the homodyne detections are parametrized by $\vec{\theta}=(\theta_{1},\theta_{2},\ldots,\theta_{M-N})$. These measurements influence the entanglement structure of the remaining modes in spatial mode $B$, effectively tailoring the time-domain entangled state into a new $N$-mode graph state,
$|\Phi_{\text{out}}\rangle=\hat{U}_{V}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})\hat{S}(\vec{r}_{\text{eff}})|0\rangle^{\otimes N},$ (1)
@@
where $\hat{U}_{V}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})$ and $\hat{S}(\vec{r}_{\text{eff}})$ denote the effective, measurement-induced operations acting on the $N$ modes. In reality, this state will also be subject to a phase-space displacement dependent on the random homodyne measurement outcomes; however, as described below, we disregard this in order to focus on the induced correlations. The parameters $\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}}$, and $\vec{r}_{\text{eff}}$ depend on the fixed circuit configuration $\{K_{i}\}$, the programmable measurement angles $\vec{\theta}$, and the initial squeezing parameters $\vec{r}_{\text{in}}$. This measurement-induced programmability enables control of the properties of the output state while maintaining the scalability and low-loss benefits of the fixed interferometer architecture. Although the long delay lines in this system introduce some loss, each optical path is traversed only once, ensuring that the total loss does not scale with the number of modes.
Thus, the resultant time-domain output of mode B mirrors the output of a traditional squeezed light interferometer, where correlations span different spatial outputs of a linear beam splitter plus phase shifter network supplied with squeezed light as in Fig. 1a. This output state is Gaussian and can therefore be described by its covariance matrix $\sigma$. To identify the effective circuit parameters for a given measurement-induced state, the covariance matrix is decomposed into a symplectic transformation $S$ of a thermal state represented by a diagonal matrix $D$ of its thermal eigenvalues (Williamson decomposition), $\sigma=\frac{1}{2}SDS^{T}$. Next, the Bloch-Messiah-Euler decomposition dissects the symplectic transformation into initial squeezing transformations $\hat{S}(\vec{r}_{\text{eff}})$ (with diagonal matrix $S_{q}$) and a passive linear transformation given by an orthogonal and symplectic matrix $L$. The unitary matrix $V$ and a corresponding unitary operator $\hat{U}_{V}(\vec{T}_{\text{eff}},\vec{\phi}_{\text{eff}})$ is extracted from $L$ (see Supplementary Material, sec. 3 [supp]). Further, the (non-unique) effective circuit parameters $\vec{T}_{\text{eff}}$, $\vec{\phi}_{\text{eff}}$ can be extracted from $V$ [Reck1994, Clements2016]. However, as mentioned above, a by-product of the homodyne measurements used to tailor the resulting Gaussian states are displacements on the remaining modes which depend on the random measurement outcomes. These must be either actively corrected by measurement feed-forward or corrected for in post-processing. If not, the effect of the accumulation of these displacements through multiple realisations will show up in the thermal eigenvalue matrix $D$. In our experimental results, we treat these displacements in post-processing at the ensemble level by setting $D=\frac{1}{2}\mathbb{I}$, thereby effectively nulling the displacements (see Supplementary Material, sec. 3 [supp]).
III Expressibility
Our scheme is not universal; it cannot reproduce an arbitrary $N$-mode squeezed light interferometer. Nevertheless, it covers a broad subset of transformations. This becomes clear when comparing the $O(N^{2})$ scaling of the number of circuit parameters in a fully programmable interferometer with the linear scaling of the programmable measurement parameters $\vec{\theta}$. Despite this restriction, the approach remains expressive enough to realize a wide range of tunable multimode transformations within a scalable and low-loss platform. To gauge the scheme’s versatility, we turn to the concept of expressibility. The conventional definition of expressibility [PQC1] compares the distribution of fidelities between pairs of output states randomly sampled from the ensemble $\mathcal{P}$ of all parametrized circuits with the equivalent distribution of states from a Haar-random reference ensemble $\mathcal{H}$. Formally, it can be quantified as the squared Hilbert-Schmidt norm $\hat{A}_{t}||_{\text{HS}}^{2}$ of the deviation from a state $t$-design,
$\hat{A}_{t}=\int_{\mathcal{H}}(\left|\psi\right\rangle\!\left\langle\psi\right|)^{\otimes t}d\psi-\int_{\mathcal{P}}(\left|\phi\right\rangle\!\left\langle\phi\right|)^{\otimes t}d\phi\ .$ (2)
As shown in the Supplementary Material, sec. 2 [supp], this is given by
$\hat{A}_{t}||_{\text{HS}}^{2}=\int_{\mathcal{H}}\!\int_{\mathcal{H}}|\langle\psi|\psi^{\prime}\rangle|^{2t}d\psi d\psi^{\prime}+\int_{\mathcal{P}}\!\int_{\mathcal{P}}|\langle\phi|\phi^{\prime}\rangle|^{2t}d\phi d\phi^{\prime}-2\int_{\mathcal{P}}\!\int_{\mathcal{H}}|\langle\psi|\phi\rangle|^{2t}d\psi d\phi\ .$ (3)
The first two terms are the $t^{\text{th}}$ moments of the intra-set fidelities over $\mathcal{H}$ and $\mathcal{P}$ and the last term is the $t^{\text{th}}$ moment of the inter-set fidelities. Intuitively, this expression can be seen to represent the difference between the distributions of the intra-set and inter-set fidelities in terms of their $t^{\text{th}}$ order moments. In practice, these expressions are estimated through sampling of states from both ensembles. The closer $\hat{A}_{t}||_{\text{HS}}^{2}$ is to zero for each $t$, the better the given parametrized circuit can reproduce the states of the reference ensemble, i.e. the more expressible it is. However, this is not immediately a useful definition for infinite-dimensional systems such as the one we consider here, where the Haar measure does not exist, making the integrals over $\mathcal{H}$ ill-defined. Another way to think of it is that the infinitely many possible states of arbitrary energy in $\mathcal{H}$ results in vanishingly few non-zero pairwise fidelities. To get a practical measure of the expressibility of our circuits, we need to tweak the definition by compactifying the space of reference states: Instead of the non-existing uniform distribution over all possible states $|\psi\rangle$, we narrow the reference ensemble down to states that can be represented similarly to the parametrized states, i.e. as a linear transformation $\hat{U}_{W}$ of squeezed input states $\hat{S}(\vec{r}_{\text{ref}})\left|0\right\rangle$. Since $W$ belongs to the compact group of unitary matrices, this can be sampled uniformly, thereby providing the most general reference ensemble for the interferometer itself. The squeezing values now need to be confined, and the most natural choice is to let them follow the same distribution as that of the measurement-induced squeezing values, $p(\vec{r}_{\text{ref}})=p(\vec{r}_{\text{eff}})$. This recipe for populating the reference ensemble $\mathcal{H}$ results in states of same average energy as the states of $\mathcal{P}$, but where the measurement-induced linear optical network $V$, which is the main object of our interest, can be compared to truly Haar-random reference networks $W$.
Figure 2: Simulated relative deviation of measurement-induced unitaries from reference unitaries. Comparison between two different measurement strategies on an initial 1D cluster state, as illustrated at the top: a straight-forward “linear” measurement pattern (a), vs. a “Knight’s distance” (KD) pattern (b). The panels show, for interferometers of sizes $N=8$, $16$ and $32$, the distributions (c-e) and the deviations (f-h) of a set of parametrized states from a random ensemble for different values of $t$ and the average photon number per mode in the output state, $\overline{n}_{pm}$. The inset provides a zoomed-in view of the datapoints corresponding to the KD strategy. Each dataset is generated by uniformly sampling the circuit parameters, with 1000 samples drawn per case, resulting in 1000 parametrized states and 1000 unitaries. The KD approach shows a clear advantage over the linear strategy, though this advantage gradually diminishes as $N$ increases.